In this equation: 3 √x -5=3 what does x equal?
Answer by Alan3354(69443)

In this equation: 3 √x -5=3 what does x equal?
------------
3*sqrt(x) -5 =3
3*sqrt(x) = 8
sqrt(x) = 8/3
x = +sqrt(8/3), -sqrt(8/3)
x = +/- (2/3)*sqrt(6)